OpenStudy (anonymous):

The population of a small town is 54,922 and has been continuously declining at a rate of 3.6% each year. What is the approximate population of the town in 10 years? 38,064 35,150 38,318 78,721

OpenStudy (anonymous):

y = ab^x ????

OpenStudy (sumi29):

Its 38,318, found using the basic compounding formula.

OpenStudy (anonymous):

\[A = P(1-r/100)^{n} = 54922 (1-3.6/100)^{10}\] answer will be a 38064

OpenStudy (sumi29):

You could do it that way too, I find it funny that two similar questions (this one and a different one by the same poster) have correct answers found using different equations. Though I think, the answer by @digitalmonk is correct.

OpenStudy (anonymous):

No, you were correct it was 38,318